what is standard form of the equation of the line passing through the point (-4,-2) and perpendicular to the line y=-2/7x+1
Answer by mananth(16946)

1 y = -2/7 x + 1
Divide by 1
y = -2/7 x + 1
Compare this equation with y=mx+b
slope m = -2/7

The slope of a line perpendicular to the above line will be the negative reciprocal
m1*m2=-1
The slope of the required line will be 7/2

m= 7/2 ,point ( -4 , -2 )
Find b by plugging the values of m & the point in
y=mx+b
-2 = -14.00 + b
b= 12
m= 7/2
The required equation is y =7/2 x + 12
m.ananth@hotmail.ca